# Provably Robust Multi-bit Watermarking for AI-generated Text

### v3

@@ Abstract
In this work, we introduce a new watermarking method for LLM-generated text grounded in pseudo-random segment assignment. We also propose multiple techniques to further enhance the robustness of our watermarking algorithm. We conduct extensive experiments to evaluate our method. Our experimental results show that our method substantially outperforms existing baselines in both accuracy and robustness on benchmark datasets. For instance, when embedding a message of length 20 into a 200-token generated text, our method achieves a match rate of $97.6\%$, while the state-of-the-art work Yoo et al. only achieves $49.2\%$. Additionally, we prove that our watermark can tolerate edits within an edit distance of 17 on average for each paragraph under the same setting.
@@ 1 Introduction
Limitations of existing text watermarking methods. Most of the existing studies in text watermarking [29, 10, 33, 34, 70, 51, 16] can only be used to detect whether a text is generated by an LLM or not. However, they are not suitable for use under the previously mentioned scenario of content source tracing. This stems from the fact that they only embed a single indicator that the text is watermarked. In response, several pioneering works [17, 63] have explored the design of multi-bit watermarking schemes for LLMs. Nevertheless, their deployments in real-world scenarios are hindered by the high computational cost of their extraction function. For example, [17] requires about 8 hours to extract a 32-bit message from a given text. In other words, these approaches lack scalability when dealing with long messages, posing significant challenges for practical deployments. Although  [68] enhances the efficiency of extraction, their method falls short in the accuracy of extracting the embedded message from watermarked texts. For example, when the message length is 16 bits, their method can only extract the embedded message from the watermarked text with a probability of 73.6%. In essence, a scalable text watermarking method that can accurately and efficiently embed and extract multi-bit messages from generated text remains elusive.
@@
To achieve an even distribution of information redundancy without losing efficiency, we propose pseudo-random segment assignment as our algorithm’s core component. We partition the entire message into multiple bit segments for embedding. During each token’s generation, we pseudo-randomly assign a segment to the token to embed. By packing multiple bits into a segment before embedding, our design makes the information redundancy distributed to each bit more balanced while avoiding the overhead of enumerating all possible messages during extraction. Our approach combines the strengths of two state-of-the-art multi-bit watermarking techniques: the high correctness of message enumeration-based methods [63, 17] and the efficiency of bit assignment-based methods [68].
We design two techniques to refine our method:
@@
  • We propose a new Large Language Model (LLM) multi-bit watermarking scheme. Our innovation lies in the design of pseudo-random segment assignment and further improvement techniques. Our scheme demonstrates a substantial enhancement in both the correctness and robustness of the multi-bit watermarking algorithm while preserving text quality.
@@
  • We validate the effectiveness of our proposed scheme through extensive experiments. Our results demonstrate a significant outperformance compared to existing multi-bit watermarking schemes, particularly in terms of the correctness of extraction, robustness against different types of attacks, and efficiency of extraction. The results also provide evidence of the practical applicability of our approach in real-world settings. We release our source code at https://github.com/randomizedtree/segment-watermark.
@@ 2.1 Zero-bit watermarking
Existing works on zero-bit watermarking. Various research has been conducted on zero-bit watermarking in text. Early approaches are mainly rule-based, such as paraphrasing [5] and synonym substitution [61]. Later, advancements in modern language models led to improved methods. In [1], they designed a watermarking framework in which both embedding and extraction are handled by text-to-text adversarially trained language models. [24, 23] embeds watermarks by context-aware lexical substitution. Recently, Kirchenbauer et al. [29] proposed imperceptible watermarks by modifying the output logits of language models during token generation. This approach has emerged as a promising approach for distinguishing language model-generated text from human-written text. [10, 33] proposed distortion-free watermarking schemes which retain the original LLM’s output distribution. Zhang et al. [69] trained a message-encoding neural network for watermark injection, followed by a message-decoding neural network for watermark extraction. Zhao et al. [70] proposed a watermarking construction that offers provable robustness guarantees for editing properties. [16, 37] proposed watermarking schemes that are publicly verifiable. [57, 38] proposed semantic-based watermarks that enhance robustness against paraphrasing for zero-bit watermarking.
@@ 2.2 Multi-bit watermarking
Multiple works [63, 17, 68] generalized the idea of selecting green tokens and adding bias to them in [29]. Wang et al. [63] proposed a multi-bit watermarking technique by using the message content as the seed to randomly select green list words. Fernandez et al. [17] consolidated multi-bit watermarks for LLMs by adopting more robust statistical tests. Concurrent with our work, Boroujeny et al. [6] extended the distortion-free watermark in [33] to the multi-bit setting. Another concurrent work Cohen et al [12] proposed a block-based multi-bit watermarking method. More similar to our work, Yoo et al. [68] proposed to pseudo-randomly assign each token a bit position to embed. Next, we provide a detailed discussion on the limitations of the existing multi-bit watermarking schemes. We also discuss the differences between our work and other works in LLM watermarking which also leverage error-correction code.
@@
Limitations of bit assignment-based multi-bit watermark. The bit assignment-based multi-bit watermark [68] pseudo-randomly assigns each token a bit position for embedding. Specifically, the bit position for the $i$-th token is pseudo-randomly determined by the $(i-1)$-th token. This approach offers high efficiency. However, a significant drawback of this method is its low correctness. The number of tokens used to embed each bit can be quite uneven due to imbalanced token frequencies. As a result, some bit positions may be represented by very few tokens or even none, making it likely to extract incorrect results for those positions.
Limitations of distortion-free multi-bit watermark. The watermark method in  [6] achieves the interesting theoretical property of exactly preserving the output distribution of the original LLM. However, our analysis reveals its poor performance when embedding longer messages in practice. This issue arises from its distribution interval encoding strategy. Specifically, the probability of generating the same token for neighboring messages is $(1-2^{1-b})^{17}$, where $b$ is the embeded message length. This probability is nearly 1 for relatively large $b$. For example, when embedding 20 bits into 200 tokens, the watermarked text will be exactly same for two neighboring messages with a probability greater than 99%. This makes it almost impossible to distinguish between the two neighboring messages. For more details, please refer to Appendix E.
@@ 4 Methodology
Table 1: Example of our watermarking and extraction results. The content in the "Prompt" column comes from datasets. The "Real Completion" column represents the ground truth completion from datasets. The "Watermarked Text" column represents text embedded with 20-bit information by our watermarking algorithm. For more examples, see Table 5 in Appendix B.
@@ 4.1.1 Insights from previous works
The algorithm in [68] is different in that it assigns each token a bit position to embed based on the hash of the previous token. In this way, tokens are divided in a pseudo-random manner. Due to the pseudo-randomness in the bit embedded by each token, this method achieves better robustness against attacks such as sentence removal. However, this method fails to achieve high accuracy when embedding longer messages. Deeper inspection into the bits extracted incorrectly reveals that these bits are often allocated few tokens or even none at all. This imbalance in token allocation significantly hinders [68] to achieve high correctness, because when a bit is embedded using only a few tokens, there is a considerable probability of extracting the bit incorrectly. In the extreme case, when no tokens are assigned to a bit, the bit’s value can only be guessed randomly.
@@ 4.1.2 Key ideas of our watermark design
Our solution is to pack multiple bits together into bit segments. During embedding, each token is pseudo-randomly assigned a segment instead of a single bit. This approach effectively reduces the number of token groups while still embedding the same message. Note that each token now simultaneously embeds every bit in its assigned segment. Consequently, the number of tokens embedding each bit is much more balanced.
@@
Our proposed approach is much more efficient during extraction compared with  [63, 17].  [63, 17] have extraction complexity $O(2^{b})$ due to the need to enumerate over the whole potential message space. In our approach, the extraction complexity is only $O(k\cdot 2^{\frac{b}{k}})$, as the enumeration is conducted for each $\frac{b}{k}$-bits segment individually.
@@ 4.1.3 Further improvements
4.2 Design details of multi-bit watermarking
Step I: In the first step, we aim to select an ECC scheme that has enough payload and satisfies the minimal code rate $R_{c}$ and minimal recover rate $R_{r}$. We first filter for all schemes $(n,k,t)_{2^{m}}\in\mathcal{F}$ satisfying $km=b$, $\frac{k}{n}\geq R_{c}$ and $\frac{t}{n}\geq R_{r}$. Among these candidate schemes, we choose the code $(n,k,t)$ with the lowest $n$ to embed. For simplicity, we use $(\hat{n},\hat{k},\hat{t})_{2^{\hat{m}}}$ to denote the selected code scheme. Algorithm 3 in Appendix C presents the complete ECC scheme selection process.
@@ 4.3 Theoretical robustness analysis
In this section, we demonstrate the robustness of our watermarking scheme against editing attempts (insertion, deletion, or substitution of a token). This represents the first LLM multi-bit watermarking scheme that provides a non-trivial provable robustness bound under edit distance [44] for every generated sentence. Some other works [70, 20, 12] also studied the provable robustness of LLM watermarking. However,  [70, 20] only apply to zero-bit watermarking. Concurrent work Cohen et al. [12] proved that if a sufficient number of entropy-based blocks remain after editing, a $1-\delta$ ratio of bits in the message can be recovered. However, it is challenging to bound how many entropy-based blocks remain after editting within some edit distance bound. Additionally, their guarantee is less compelling because it cannot prove perfect extraction.
The edit distance between two strings is defined as the minimum number of operations (insertion, deletion, or substitution of a token) required to transform one string into the other. Before presenting the formal theorem, it is essential to first introduce one notation.
@@
Through the theorem, we demonstrate that when the attacker modifies the watermarked text within an edit distance budget, with quite a small probability, the extracted information will be manipulated. For example, when $T=200$, $\eta=15$, $b=20$, $k=5$, $m=4$, $n=6$, $\textbf{c}=[30,35,35,30,35,35],\leavevmode\nobreak\ \textbf{d}=[25,31,31,26,32,30]$, the upper bound of extraction error rate is $0.08\%$. Therefore, it is almost impossible for the attacker to distort our watermark within this $\eta\leq 15$ budget. Due to the space limitation, the proof is available in Appendix A.
@@
The robust bound $B$ can be determined by performing a binary search to find the maximum value of $\eta$. The algorithm for computing the robust bound $B$ for a specific paragraph is detailed in Algorithm 5 in Appendix C. The complexity of Algorithm 5 is $O(n\cdot\eta_{max}^{4}+t\cdot\log\eta_{max})$, dominated by the computation of the $\mathsf{PI}$ array. In practice, we typically use $n\leq 10$ and $\eta_{max}=40$, allowing $B$ to be efficiently computed within a few seconds on a modern CPU.
4.4 Detect machine-generated text
In real-world applications, such as content auditing, it is often crucial to initially determine whether the text is watermarked and subsequently extract the message embedded in the watermarked text. It would be meaningless to extract information from natural text that does not even carry a watermark.
Our method can also detect watermarked text from human-written text in a manner similar to [30]. The intuition is that our multi-bit watermarking method also injects token bias into generated text. This allows us to leverage the statistical difference between natural text and machine-generated text for detection purposes.
Although the detector has no knowledge about the message embedded in the watermark, the $\mathsf{COUNT}$ matrix is still computable. Due to added token bias, for segment $p$ it is quite likely that the segment value that has the most green tokens is the correct value. Therefore, we can approximate the total number of green tokens in the paragraph by using the sum of the maximum green token numbers for each segment. From Figure 3 we can see that all of the blue dots are close to the red line $y=x$. This indicates that the sum of max statistics we use is very close to the true number of green tokens. Algorithm 4 in Appendix C gives a precise description of our machine text detection algorithm.
Figure 3: Green token number and sum of maximum number of green tokens for different watermarked paragraphs. The two statistics for machine text detection are similar showed in the figure.

### v5

@@ Abstract
In this work, we introduce a new watermarking method for LLM-generated text grounded in pseudo-random segment assignment. We also propose multiple techniques to further enhance the robustness of our watermarking algorithm. We conduct extensive experiments to evaluate our method. Our experimental results show that our method achieves a much better tradeoff between extraction accuracy and time complexity, compared with existing baselines. For instance, when embedding a message of length 20 into a 200-token generated text, our method achieves a match rate of $97.6\%$, while the state-of-the-art work Yoo et al. only achieves $49.2\%$. Additionally, we prove that our watermark can tolerate edits within an edit distance of 17 on average for each paragraph under the same setting.
@@ 1 Introduction
Limitations of existing text watermarking methods. Most of the existing studies in text watermarking [33, 11, 36, 37, 74, 53, 18, 16, 68, 22, 29, 46] can only be used to detect whether a text is generated by an LLM or not. However, they are not suitable for use under the previously mentioned scenario of content source tracing. This stems from the fact that they only embed a single indicator that the text is watermarked. In response, several pioneering works [19, 65] have explored the design of multi-bit watermarking schemes for LLMs. Nevertheless, their deployments in real-world scenarios are hindered by the high computational cost of their extraction function. For example, [19] requires about 8 hours to extract a 32-bit message from a given text. In other words, these approaches lack scalability when dealing with long messages, posing significant challenges for practical deployments.  [71] proposed to pseudo-randomly allocate short bit segments (1-3 bits) to enhance extraction efficiency. However, their method falls short in the accuracy of extracting the embedded message from watermarked texts. For example, when the message length is 16 bits, their method can only extract the embedded message from the watermarked text with a probability of 73.6%. In essence, a scalable text watermarking method that can accurately and efficiently embed and extract multi-bit messages from generated text remains elusive.
@@
To achieve an even distribution of information redundancy without losing efficiency, we design an approach unifying both message enumeration-based methods and bit assignment-based methods. We partition the entire message into multiple bit segments for embedding. During each token’s generation, we pseudo-randomly assign a segment to the token and embed it leveraging ideas in enumeration-based watermarks. Our design makes the information redundancy distributed to each bit more balanced while avoiding the overhead of enumerating all possible messages during extraction. Our method offers a systematic way to balance the trade-off between watermark accuracy and computational efficiency.
We also design two techniques to refine our method:
@@
  • We propose a new Large Language Model (LLM) multi-bit watermarking scheme. Our innovation lies in the watermark design which unifies message enumeration-based methods and bit assignment-based methods. We also proposed further improvement techniques. Our scheme offers a systematic approach to balancing the trade-off between watermark accuracy and computation efficiency in multi-bit watermarking algorithm, while simultaneously enhancing the watermark robustness and preserving text quality.
@@
  • We validate the effectiveness of our proposed scheme through extensive experiments. Our results demonstrate a significant outperformance compared to existing multi-bit watermarking schemes, particularly in terms of the correctness of extraction, robustness against different types of attacks, and efficiency of extraction. We release our source code at https://github.com/randomizedtree/segment-watermark.
@@ 2.1 Zero-bit watermarking
Existing works on zero-bit watermarking. Various research has been conducted on zero-bit watermarking in text. Early approaches are mainly rule-based, such as paraphrasing [5] and synonym substitution [63]. Later, advancements in modern language models led to improved methods. In [1], they designed a watermarking framework in which both embedding and extraction are handled by text-to-text adversarially trained language models. [27, 26] embeds watermarks by context-aware lexical substitution. Recently, Kirchenbauer et al. [33] proposed imperceptible watermarks by modifying the output logits of language models during token generation. This approach has emerged as a promising approach for distinguishing language model-generated text from human-written text.  [11, 36, 68] proposed distortion-free watermarking schemes which retain the original LLM’s output distribution. Zhang et al. [73] trained a message-encoding neural network for watermark injection, followed by a message-decoding neural network for watermark extraction. Zhao et al. [74] proposed a watermarking construction that offers provable robustness guarantees for editing properties. [59, 40] proposed semantic-based watermarks that enhance robustness against paraphrasing for zero-bit watermarking. [22] formulated watermarking as a search problem that aims to identify the token chunks with the lowest p-value among the top outputs of LLM.
@@ 2.2 Multi-bit watermarking
Multiple works [65, 19, 71] generalized the idea of selecting green tokens and adding bias to them in [33]. Wang et al. [65] proposed a multi-bit watermarking technique by using the message content as the seed to randomly select green list words. Fernandez et al. [19] consolidated multi-bit watermarks for LLMs by adopting more robust statistical tests. Concurrent with our work, Boroujeny et al. [7] extended the distortion-free watermark in [36] to the multi-bit setting , while [72] generalized [11] to achieve multi-bit distortion-free watermarking. Another concurrent work Cohen et al [13] proposed a block-based multi-bit watermarking method. More similar to our work, Yoo et al. [71] proposed to pseudo-randomly assign each token a short bit segment (1-3 bits) to embed. Next, we provide a detailed discussion on the limitations of the existing multi-bit watermarking schemes. We also discuss the differences between our work and other works in LLM watermarking which also leverage error-correction code.
@@
Limitations of bit assignment-based multi-bit watermark. The bit assignment-based multi-bit watermark [71] pseudo-randomly assigns each token a short bit-segment (1-3 bits) for embedding. Specifically, the segment for the $i$-th token is pseudo-randomly determined by the $(i-1)$-th token. This approach offers high efficiency. However, a significant drawback of this method is its low correctness. The number of tokens used to embed each bit can be quite uneven due to imbalanced token frequencies. As a result, some bit positions may be represented by very few tokens or even none, making it likely to extract incorrect results for those bits.
Although the imbalance issue of  [71] could be mitigated by increasing the segment size, this results in adding bias on a very small vocabulary partition, which significantly degrades text quality.
Limitations of distortion-free multi-bit watermark. The watermark method in  [7] achieves the interesting theoretical property of exactly preserving the output distribution of the original LLM. However, our analysis reveals its poor performance when embedding longer messages in practice. This issue arises from its distribution interval encoding strategy. Specifically, the probability of generating the same token for neighboring messages is $(1-2^{1-b})^{17}$, where $b$ is the embeded message length. This probability is nearly 1 for relatively large $b$. For example, when embedding 20 bits into 200 tokens, the watermarked text will be exactly same for two neighboring messages with a probability greater than 99%. This makes it almost impossible to distinguish between the two neighboring messages. For more details, please refer to Appendix C.
[72] extended the approach of [11] to enable distortion-free multi-bit watermarking by additionally feeding the $i$-th bit of the message to the pseudo-random function for the $i$-th token. However, this method is highly vulnerable to attacks like sentence removal.
@@ 4 Methodology
Table 1: Example of our watermarking and extraction results. The content in the "Prompt" column comes from datasets. The "Real Completion" column represents the ground truth completion from datasets. The "Watermarked Text" column represents text embedded with 20-bit information by our watermarking algorithm.
@@ 4.1.1 Insights from previous works
The algorithm in [71] is different in that under typical setting, it assigns each token a 2 bit segment of the message to embed based on the hash of the previous token. In this way, tokens are divided in a pseudo-random manner. Due to the pseudo-randomness in the bit segment embedded by each token, this method achieves better robustness against attacks such as sentence removal. However, this method fails to achieve high accuracy when embedding longer messages. Deeper inspection into the bits extracted incorrectly reveals that these bits are often allocated few tokens or even none at all. This imbalance in token allocation significantly hinders [71] to achieve high correctness, because when a segment is embedded using only a few tokens, there is a considerable probability of extracting the segment incorrectly. In the extreme case, when no tokens are assigned to a segment, the segment’s value can only be guessed randomly.
@@ 4.1.2 Key ideas of our watermark design
Our solution is to pack multiple bits together into bit segments. The difference between our approach and [71] lies in the method of embedding a segment into a token. [71] embeds a $d-$bit segment by adding bias to a $\frac{1}{2^{d}}$ fraction of vocabulary. Increasing the segment size of [71] (e.g., size 8) could mitigate the assignment imbalance but will split the vocabulary into 256 parts per step. This approach severely restricts word choice for each token, leading to a decline in text quality. Due to this limitation, rather than using longer bit segments, [71] chose to assign a 2-bit segment to each token. Contrarily, our solution embeds the assigned multi-bit segment into a token by incorporating it as an additional input to the hash function. Our approach effectively reduces the number of segments compared with [71], while still embedding the same message. Note that each token now simultaneously embeds every bit in its assigned segment. Consequently, the number of tokens embedding each bit is much more balanced.
@@
Our proposed approach is much more efficient during extraction compared with  [65, 19].  [65, 19] have extraction complexity $O(2^{b})$ due to the need to enumerate over the whole potential message space. In our approach, the extraction complexity is only $O(k\cdot 2^{\frac{b}{k}})$, as the enumeration is conducted for each $\frac{b}{k}$-bits segment individually. Our method provides a systematic way to trade-off between watermark accuracy and computational efficiency by using different segment number $k$. A smaller $k$ improves watermark accuracy by allowing more tokens to embed each segment but increases computational cost. Conversely, a larger $k$ reduces extraction complexity but also slightly sacrifices the watermark accuracy.
@@ 4.1.3 Further improvements
4.2 Design details of multi-bit watermarking
Step I: In the first step, we aim to select an ECC scheme that has enough payload and satisfies the minimal code rate $R_{c}$ and minimal recover rate $R_{r}$. We first filter for all schemes $(n,k,t)_{2^{m}}\in\mathcal{F}$ satisfying $km=b$, $\frac{k}{n}\geq R_{c}$ and $\frac{t}{n}\geq R_{r}$. Among these candidate schemes, we choose the code $(n,k,t)$ with the lowest $n$ to embed. For simplicity, we use $(\hat{n},\hat{k},\hat{t})_{2^{\hat{m}}}$ to denote the selected code scheme. Algorithm 3 in Appendix B presents the complete ECC scheme selection process.
@@ 4.3 Theoretical robustness analysis
In this section, we demonstrate the robustness of our watermarking scheme against editing attempts (insertion, deletion, or substitution of a token). This represents the first LLM multi-bit watermarking scheme that provides a non-trivial provable robustness bound under edit distance [47] for every generated sentence. Before presenting the formal theorem, it is essential to first introduce one notation.
@@
Through the theorem, we demonstrate that when the attacker modifies the watermarked text within an edit distance budget, with quite a small probability, the extracted information will be manipulated. Due to the space limitation, the proof of the theorem is available in Appendix A.
@@
The robust bound $B$ can be determined by performing a binary search to find the maximum value of $\eta$. The algorithm for computing the robust bound $B$ for a specific paragraph is detailed in Algorithm 4 in Appendix B. The complexity of Algorithm 4 is $O(n\cdot\eta_{max}^{4}+t\cdot\log\eta_{max})$, dominated by the computation of the $\mathsf{PI}$ array. In practice, we typically use $n\leq 10$ and $\eta_{max}=40$, allowing $B$ to be efficiently computed within a few seconds on a modern CPU.
4.4 Security analysis
In cryptography, hash functions are often modeled under the random oracle model (ROM) [20, 6]. This model is widely deployed for analysis in cryptography. The random oracle $H:\{0,1\}^{*}\rightarrow\{0,1\}^{n}$ is defined as follows:
  • For every input $x,H(x)$ produces a uniformly random $n$-bit output.
  • The response is the same for identical queries, but is independent with other inputs.
We first analyze the theoretical security of our watermark’s secret key under the random oracle model (ROM), which means that $\mathsf{hash}(sk,\cdot)$ is viewed as a random oracle. The output of the hash function cannot be directly observed by the watermark’s secret key extraction adversary. To simplify analysis, we consider a stronger adversary $\mathcal{A^{*}}$, formalized as:
  • $\mathcal{A^{*}}$ observes a set of pairs $\left\{\left(X_{j},\mathsf{hash}\left(\mathrm{sk},X_{j}\right)\right)\right\}_{j=1}^{q}$, where $X_{j}$ are adaptively chosen bit-strings, and $q$ is the number of queries.
  • After $q$ queries, $\mathcal{A^{*}}$ outputs a candidate for $sk$.
Then, the problem becomes analyzing whether it is possible to extract the secret key by observing hash outputs. Due to computation power limitation, assume $\mathcal{A^{*}}$ tests at most $Q$ key guesses, $\{sk^{\prime}_{1},\ldots,sk^{\prime}_{Q}\}$. The test is done by checking $\mathsf{hash}(sk^{\prime}_{i},X_{j})=\mathsf{hash}(sk,X_{j}),j=1,\ldots,q$. Due to the uniform randomness of ROM, the advantage of $\mathcal{A^{*}}$ of finding the exact $sk$ in $Q$ queries compared with completely random guessing is bounded by:
$\operatorname{Pr}[\mathcal{A^{*}}\text{ succeeds }]-\operatorname{Pr}[\mathcal{A^{*}}\text{ succeeds by random guessing}]\leq\frac{Q}{2^{n}}$
In contrast, if $\exists$ adversary $\mathcal{A^{*}}$ recover $sk$ with probability advantage significantly higher than $\frac{Q}{2^{n}}$, it indicates that $\mathcal{A^{*}}$ already leverages the structure of the output of $\mathsf{hash}$. This breaks the assumption that $\mathsf{hash}$ (viewed as a random oracle) provides independent random outputs.
In practice, the output bit length of the hash function is usually $n\geq 128$. The computation power of $\mathcal{A^{*}}$ is assumed as $Q\leq 2^{64}$. Thus the probability of $\mathcal{A^{*}}$ to succeed in extracting the secret key is no bigger than $\frac{Q+1}{2^{n}}\approx 2^{-64}$, which is negligible.
From the above analysis, we can see that it is very difficult to extract the secret key even by directly observing the hash outputs. Thus we prove that it is very difficult for an adversary to recover the secret key by observing watermarked text.
Next, we consider the security of our watermark against spoofing attacks. In the case of spoofing attack, the adversary has access to $k$ ($k\geq 1$) distinct accounts, each associated with a unique user ID $m_{1},m_{2},...,m_{k}$. For simplicity of analysis, we assume only the userID is embedded. The adversary’s objective is to forge a piece of text that will be considered as originating from the target user ID $m_{t}$ ($m_{t}\notin\{m_{1},...,m_{k}\}$).
Similarly, we also consider a stronger adversary $\mathcal{A^{\prime}}$, who can adaptively query the hash function with user IDs among $m_{1},...,m_{k}$ and observe the output. The target of the adversary is to guess the hash output of $\mathsf{hash}(sk,m_{t})$ based on observations.
Based on the random oracle assumption, the hash results of different inputs are uniformly random. Therefore, $\forall m_{t}\neq m_{i}(i\in\{1,2,...,k\})$, $\mathsf{hash}(sk,m_{t})$ and $\mathsf{hash}(sk,m_{i})$ are independent random variables. In other words, the adversary $\mathsf{A^{\prime}}$ gain no advantage in predicting $\mathsf{hash}(sk,m_{t})$ by querying $\mathsf{hash}(sk,m_{i})$. Consequently, the hash value $\mathsf{hash}(sk,m_{t})$ remains uniformly random from the adversary’s perspective. As the actual adversary in spoofing attack is strictly weaker than the adversary $\mathcal{A^{\prime}}$ (he cannot observe any actual hash values), our watermark is theoretically resistant to spoofing attack under ROM.